# Ultrasonic Medical Tissue Imaging Using Probabilistic Inversion: Leveraging Variational Inference for Speed Reconstruction and Uncertainty Quantification

### v2

@@ 7 Conclusion
In this study, an SVGD-based FWI algorithm is employed for medical ultrasound imaging, demonstrating its effectiveness in reconstructing the SOS distributions of imaged objects. For comparison, the conventional FWI and SVI are also used to reconstruct the SOS. The results show that all three algorithms provide accurate estimates of object properties, with SVGD offering additional benefits, such as faster convergence and the ability to quantify uncertainty. The SVGD results exhibit low relative errors and minimal uncertainty in the core regions of the object. Furthermore, the performance of three algorithms is evaluated using a realistic breast tissue model. The SVGD algorithm demonstrates comparable or even superior accuracy to other two algorithms. The uncertainty quantified by SVGD aligns more closely with the true SOS variations and provides tighter bounds for the prior distribution, positioning SVGD-based FWI as a promising tool for inversion imaging tasks.
@@
8 Acknowledgement
This work was supported by the National Natural Science Foundation of China (grant numbers 12122403, 12034005, and 12327807). The computations for this research were performed using the CFFF platform at Fudan University. Thanks to Oscar Bates for his constructive suggestions.
9 Data availability
Data will be made available on reasonable request.

### v3

@@ 4 Discussion
A practical limitation of the SVGD-based FWI framework lies in its computational cost, particularly for large-scale inverse problems. Each particle requires one forward and one adjoint wave equation solve per iteration, resulting in a total of $2n$ wave equation solves. Although this increases the computational burden compared to deterministic FWI, empirical results (e.g., Figures 5 and 9) demonstrate that satisfactory inversion accuracy and uncertainty quantification can be achieved with as few as 5–10 particles. This flexibility offers a trade-off between computational efficiency and the fidelity of the posterior distribution. Additionally, since the forward and adjoint wave equation solves for different particles are independent, the method is inherently parallelizable. By leveraging high-performance computing resources and executing computations in parallel across particles, the wall-clock time can be significantly reduced, making SVGD-based inversion feasible for more complex and computationally demanding imaging tasks.
@@ 5 Conclusion
In this study, an SVGD-based FWI algorithm is employed for medical ultrasound imaging, demonstrating its effectiveness in reconstructing the SOS distributions of imaged objects. For comparison, the conventional FWI and SVI are also used to reconstruct the SOS. The results show that all three algorithms provide accurate estimates of object properties, with SVGD offering additional benefits, such as faster convergence and the ability to quantify uncertainty. The SVGD results exhibit low relative errors and reliable uncertainty in the core regions of the object. Furthermore, the performance of three algorithms is evaluated using a realistic breast tissue model. The SVGD algorithm demonstrates comparable or even superior accuracy to other two algorithms. Moreover, the uncertainty quantified by SVGD aligns closely with the benchmark MH-MCMC method, reflecting SOS variations of the inversion imaging. Sensitivity analysis of particle count shows that even with fewer particles (e.g., $n=10$ or $15$), high-quality inversion results are achievable. When moderate uncertainty resolution is acceptable, 5 particles can provide satisfactory reconstructions, thereby enabling a balance between computational efficiency and the fidelity of uncertainty quantification. The results underscore the potential of SVGD-based FWI as a promising tool for inversion imaging tasks.
@@
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China (grant numbers 12122403, 12034005, and 12327807) and National Key R&D Program of China (2023YFC2410800). The computations for this research were performed using the CFFF platform at Fudan University. Thanks to Oscar Bates for his constructive suggestions.
7 Declaration of Competing Interest
The authors declare that there are no known competing interests that could have influenced the work reported in this paper.